OpenStudy (pphalke):

the perimeter of a basketball court is 108 meters and the length is six meters longer than twice the width. what are the length and width?

OpenStudy (parker.goodbar):

length is 30 width is 24

OpenStudy (anonymous):

Length=2x+6 Equation (Combine all 4 sides) 6x+12=108 x=16 Length=2(16)+6 Final Answer: Length=38 Width=16

OpenStudy (calculusxy):

You can write a system of equations to solve for the length and width. One of them could be the formula for finding the perimeter: \(2l + 2w = 108\) And the other one could be that the length is six meters less than twice the width: \(l = 2w - 6\). So you have: \(2l + 2w = 108\) \(l = 2w - 6\) I prefer for you to use substitution for this particular method because you already have what \(l\) is equal to and you can substitute that value into the first equation (\(2l + 2w = 108\)).

OpenStudy (parker.goodbar):

thats not right at all because the length is only 6 meters longer 38 and 16 are a whole 22 meters longer

OpenStudy (parker.goodbar):

24x2=48+(30x2)=108 on this the length is only 6 meters larger which is what the equation is asking for you to do